# Thermodynamics of the dipole-octupole pyrochlore magnet CeHfO in applied magnetic fields

### v1

@@ Abstract
The recently discovered dipole-octupole pyrochlore magnet CeHfO is a promising three-dimensional quantum spin liquid candidate which shows no signs of ordering at low temperature. The low energy effective pseudospin-1/2 description in a magnetic field is characterized by the XYZ Hamiltonian and a Zeeman term where the dipolar local z-component of the pseudospin couples to the local z-component of the applied magnetic field, while the local x- and y-components of the pseudospin remain decoupled as a consequence of their octupolar character. Using effective parameters determined in V. Poree et al., arXiv:2305.08261 (2023), remarkable experimental features can be reproduced, as for instance the specific heat and magnetization data as well as the continuum of states seen in neutron scattering. Here we investigate the thermodynamic response to magnetic fields applied along the global [110] direction using specific heat measurements and fits using numerical methods, and solve the corresponding magnetic structure using neutron diffraction. Specific heat data in moderate fields are reproduced well, however, at high fields the agreement is not satisfactory. We especially observe a two-step release of entropy, a finding that demands a review of both theory and experiment. We address it within the framework of three possible scenarios, including an analysis of the crystal field Hamiltonian not restricted to the two-dimensional single-ion doublet subspace. We conclusively rule out two of these scenarios and find qualitative agreement with a simple model of field misalignment with respect to the crystalline direction. We discuss the implications of our findings for [111] applied fields and for future experiments on CeHfO and its sister compounds.
@@ I Introduction
In pyrochlores, the magnetic ions form a lattice of tetrahedra connected by their summits, and the basic motivation came from spin ice physics, originally discovered in HoTiO and DyTiO [22, 23, 24, 25]. In these materials, a strong crystal field forces dipole magnetic moments to align along or against the local axis (local axis) directed from the site to the center of the tetrahedron. Such systems do not show any long range order down to mK temperatures, but magnetic configurations are formed by an organizing principle, the ‘2-in 2-out ice rule’, emerging from effective ferromagnetic interactions. This ice rule constraint manifests itself in various ways, notably the observation of Pauling’s residual entropy [24] and the existence of a classical Coulomb liquid with characteristic ‘pinch points’ [26] accessible in neutron scattering.
@@
One approach to find QSLs has been to look for pyrochlore magnets with relaxed constraint on the Ising anisotropy of the magnetic moments. This approach, however, has proved difficult. For example, the case of YbTiO [27, 28] was anticipated to be a QSL but was found to have a ferromagnetic ground state [29, 30, 31, 32, 33]. Alternatively, a promising route for QSL physics was found with materials based on the non-Kramers trivalent Pr ion having Ising-like dipole magnetic moments [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], and where interactions between electric quadrupoles [45] and residual non-magnetic disorder [41, 43] can both be source of quantum fluctuations to stabilize a QSL.
@@
More recently, the case of trivalent rare-earth ions that can stabilize a ‘dipole-octupole’ ground state doublet, such as Nd or Ce, has been actively investigated and appears very promising. As in most rare-earth pyrochlores, since the crystal-electric field (CEF) ground state doublet is well separated from the rest of the spectrum by a gap of few tens of meV, microscopic descriptions use an effective spin 1/2 degree of freedom. The special feature of these Kramers ‘dipole-octupole’ doublets lies in the fact that the component of the pseudo-spin essentially represents the magnetic (dipolar) moment, while the the and components are built from octupolar moments. The theoretical calculations presented in this work are based on a minimal nearest neighbor (nn) XYZ Hamiltonian incorporating all symmetry-allowed effective interactions between these pseudo-spins on the pyrochlore lattice [46] and accounting for the applied magnetic field when projected onto the ground doublet subspace:
@@
The basis is chosen such that is defined in the local direction on a given rare earth site and leads to . and are effective coupling strengths and refers to nearest neighbor pairs of sites. We note that, in the chosen basis with , the Zeeman term couples only the local component of the magnetic field to the dipolar component while the octupolar ( and ) components do not contribute to this term.
@@
Depending on the relative strength of , different situations can emerge. For instance, in NdZrO, is dominant and negative, leading to an antiferromagnetic all-in–all-out ground state, with however significant dynamics arising from a non-negligible ‘transverse’ term [47, 48, 49, 50]. In the Ce-based pyrochlore materials CeSnO, CeZrO and CeHfO [51, 52, 53, 54, 55, 56, 57], studies concluded to a dominant positive interaction between and/or pseudo-spin components, hence a new class of spin ices where the ice rule is ‘octupolar’. As a result, the magnetic diffuse scattering in CeSnO does not show the familiar characteristic signature observed in the dipolar spin ice systems. Some magnetic diffuse scattering, however, arises at high momentum transfers and is observable with neutrons thanks to the weak coupling of octupoles to neutrons [51, 56].
Furthermore, in all three Ce-based pyrochlores, a continuum of spectral weight [53, 54] – the hallmark of fractionalized excitations in a QSL, has been observed in neutron scattering in zero applied magnetic field. In CeHfO, effective exchange parameters were reported previously by some of us [57] with a dominant interaction:
@@
(all ’s in meV, dimensionless), which locates the material of the present study in the -flux U(1) QSL phase [58, 46, 59, 60, 61, 62, 63], as suggested also for CeZrO and CeSnO. (However, a recent study on samples of CeSnO has suggested the possibility of AIAO order [64]) The interaction parameters result in a rich phase diagram for cerium pyrochlores, and their tiny energy scales result in a potentially large response to small changes in the material composition. For instance, this is illustrated in the markedly different exchange parameters in CeZrO and CeHfO despite Zr and Hf sharing very similar ionic radii and electronegativities. The highly tunable nature of these tiny energy scales may also explain the contrasted magnetic neutron scattering response of CeSnO samples prepared at lower temperature [64] compared to other studies [52, 51, 52, 65]. The methodology to deduce reliable sets of exchange parameters from the available experimental data is a central problem in the study of these materials.
@@
In the case of dipolar systems, the application of external magnetic fields has been a valuable resource for building confidence in the microscopic description of such materials [27, 25]. Large fields spin polarize the system allowing direct measurement of the energy- and momentum-resolved spin waves (for example, see the case of YbTiO [27, 30, 31, 28] and references therein). Using linear spin wave theory, it is then possible to accurately determine the underlying interactions in these dipolar spin ices. The case of dipolar-octupolar magnets is starkly different [46, 59, 60, 61, 62, 63]: a field-induced magnetic order can exist, but with no spectral weight associated with the octupolar low energy excitations measured with neutrons at low momentum transfers. Absence of spin waves has also been noted in CeZrO, despite the measurement of a field-induced magnetic order in a finite magnetic field [66] These features outline the importance of methods to evaluate exchange interactions of cerium pyrochlores from measured data.
@@
Given this prelude to dipole-octupole pyrochlore magnets, in this work we present neutron diffraction, magnetization and specific heat data on CeHfO for magnetic fields applied along the crystallographic [110] direction. Strikingly, we infer a two-step release of entropy from the specific heat measurements, a finding that demands a review of the experiment and exploration of several theoretical scenarios. Using a combination of quantum many-body calculations and analytical insights, we present strong evidence for the importance of a field misalignment away from the [110] axis as the origin of the peculiar evolution of entropy as a function of the applied field strength. We also revisit results for the [111] field direction, previously addressed elsewhere [57].
@@
The rest of the paper is organized as follows. In section II we describe the setup of the pyrochlore lattice, discussing the phases expected at high magnetic fields. In section III we discuss important details of our experiments, and the results of measurements for magnetic fields, highlighting the agreement between new data for moderate [110] fields with the prediction from our previously determined model parameters given in Eq. (2). Section IV is devoted to our theoretical analysis for [110] fields, while Section V provides, for completeness, a similar analysis for the [111] fields, with simple torque calculations confirming that the effects of a misalignment are far less important in that case. Finally, in Section VI we conclude and discuss the implication of our work on future experiments.
@@ II Pyrochlore Setup
We distinguish the global coordinate system, which we denote in what follows, from the local coordinate system at each rare earth site, which is given, following Ref. [46], as follows:
@@
Equation (1) indicates that only the local components is of consequence in determining the coupling of the applied external magnetic field to a given pseudo-spin.
An interesting situation arises for a magnetic field of strength applied perfectly along [110], i.e., and in the global basis. The Zeeman term is exactly zero for sublattices 1 and 2, i.e. the spins do not directly couple to the applied [110] magnetic field, for arbitrary field strengths. The Zeeman term is non-zero for sublattices 0 and 3, i.e., the sites. For the case of large field strengths, the ground state corresponds to the pseudospin on sublattice 0 pointing “out” and the one on sublattice 3 pointing “in” i.e. along the local z axes, away and towards the center of the tetrahedron, respectively. This configuration has been qualitatively depicted in Fig. 1(b).
If the spins were truly free, the ground state would be exactly degenerate resulting in a residual entropy of per site. In practice, the spins do couple to the spins which in turn couple to the magnetic field, thus the spins are essentially free only at temperature scales above the scale of effective magnetic interactions. One should, however, expect to see the signatures of this effective decoupling in the specific heat and the associated magnetic entropy, especially at high fields where the dominant scale is set by the Zeeman term. Rather curiously, we find that our high-field experiments do not necessarily adhere to this expectation, a finding we address further in the next sections.
@@ III Experimental results
Here we present the details of the experiments performed, along with the key outcomes and puzzles they pose. In subsection III.1, we show neutron diffraction data in [110] fields, which provide a picture consistent with the spins ordering while the spins remain disordered. We then discuss in subsection III.2 specific heat results to be compared with numerical calculations presented in section IV.
@@ III.1 Magnetic neutron diffraction and magnetization
The induced magnetic structure under application of a magnetic field parallel to the direction was investigated using single crystal neutron diffraction (Zebra, SINQ, PSI). The single crystal sample of CeHfO was cooled down using a dilution refrigerator inside a 10 T vertical cryomagnet. Reflections in and out of the horizontal scattering plane ( degrees) were collected using thermal neutrons (1.18 Å) via a He point detector. The neutron beam was monochromatized and vertically focused using the (311) Bragg reflection of a germanium monochromator. Measurements took place at 1.5 K, 0.2 K and 0.05 K, under the application of a magnetic field = 0 T, 0.5 T, 2 T and 8 T. Magnetic Bragg intensities were extracted by subtraction of their respective nuclear contributions measured without magnetic field. All magnetic reflections could be accounted for using a propagation vector. Other directions of reciprocal space were investigated in order to check for magnetic Bragg scattering of a different propagation vector, which confirmed that the magnetic structure can be described within the chemical unit cell.
Fits of the experimental integrated intensities corrected for Lorentz factor were performed using the FullProf Suite program [67] assuming the formation of and chains and a 2-in-2-out local constraint (fit displayed in Fig. 2), owing to the strong anisotropy and frustration experienced by the magnetic moments [68]. Across all temperatures and non-zero magnetic fields, the determined magnetic structures revealed a net magnetic moment on the sites, lying along their respective local 111 easy axis, polarized in the field direction and increasing with field strength. No net magnetic moment on the sites could be identified for = 0 T, 0.5 T, 2 T. Interestingly, for a magnetic field strength of = 8 T, allowing a small magnetic moment (0.100.05 ) parallel to the magnetic field leads to a slight improvement of the fit. This could be due to the increased number of fit parameters (and therefore be artificial) or could alternatively be real, potentially stemming from a slight misalignment of the field with respect to the ideal direction or due to a reduced single-ion anisotropy inherited from defective CEF environments [55].
Magnetization vs field was measured using SQUID magnetometers equipped with a miniature dilution refrigerator developed at the Institut Néel-CNRS Grenoble [69]. We report data for the magnetic field applied along the [110] direction, as shown in Fig. 2(b) for data at 0.08 K and in Fig. 3(d) at four representative temperatures. The magnetization of the sample was also calculated based on the magnetic structure determined from neutron diffraction at 0.05 K and compared with the macroscopic magnetization at 0.08 K (Fig. 2(b)), showing a good agreement and further corroborating the -chain/-chain structure ansatz.
@@ III.2 Specific heat
A thin slice of CeHfO was prepared for specific heat measurement using the single crystal reported in [55] and a tungsten wire saw. The [110] orientation of the slice’s faces was assessed using a x-ray Laue diffractometer. Specific heat measurements between 0.05 K and 0.8 K were performed using a home-built calorimeter inserted in a dilution refrigerator at Université de Sherbrooke. We will refer to this as the “experimental setup A”. In this setup, the heater and the thermometer were glued directly to the sample using GE varnish. This assembly was suspended on zylon wires to decouple it thermally from the copper mount. Electrical contacts were made using 7 µm diameter superconducting NbTi wires. The thermal conductivity between the sample assembly and the copper mount was adjusted using a manganin wire, in order to have suitable relaxation times for the measurements. The manganin heat link was 1.7 mm long with a diameter of 0.1 mm. As well as providing the thermal link to the refrigerator, this short length of wire also provided significant structural support to the sample in addition to the very thin zylon filaments. The sample was oriented to apply the magnetic field along the [110] crystallographic direction. Data were acquired using the quasi-adiabatic heat pulse method.
@@
Specific heat measurements between 0.4 K and 15 K were carried out using a Quantum Design physical properties measurement system (PPMS), with vertically applied magnetic fields up to 6 T. This is labeled as the “experimental setup B” in the rest of the paper. For this setup, the sample was mounted with its smoothest side facing down towards the puck’s platform using an Apiezon grease, allowing the application of a magnetic field parallel to the [110] crystallographic direction. The platform itself is suspended by four wires, which also provide electrical connections to the heater and thermometer. Data measured on the nonmagnetic analog LaHfO were used to subtract lattice contributions.
@@
Figures 3(a),(b),(c) and 5 show experimental specific heat data for CeHfO in a magnetic field along the [110] direction in both setup A (green) and B (black), for low ( T) and high field strength respectively. Note that for the latter case, only data from setup B is available. The experimental results reveal a prominent two-step release of entropy, which appears as two outstanding bumps in the specific heat. Note that such a two-step entropy release is absent (or at least not prominent) for fields applied in the [111] direction [53]. The temperature of both bumps in the specific heat increases with the applied field strength.
@@
As emphasized above, the field applied along the [110] direction separates the spins in two types of chains, labeled and . Owing to the amplitude of the various effective interactions, the high-temperature bump seems simply to be due to the Zeeman term of the spins, scaling linearly with the field. The low temperature feature appears more interesting and seems to indicate correlations on the -chains, as investigated in more detail in the next section.

### v2

# Thermodynamics of the dipole-octupole pyrochlore magnet Ce2Hf2O7 in applied magnetic fields

@@ Abstract
The recently discovered dipole-octupole pyrochlore magnet Ce2Hf2O7 is a promising three-dimensional quantum spin liquid candidate which shows no signs of ordering at low temperature. The low energy effective pseudospin-1/2 description in a magnetic field is characterized by the XYZ Hamiltonian and a Zeeman term where the dipolar local -component of the pseudospin couples to the local -component of the applied magnetic field, while the local - and -components of the pseudospin remain decoupled as a consequence of their octupolar character. Using effective Hamiltonian parameters determined in V. Poree et al., arXiv:2305.08261 (2023), remarkable experimental features can be reproduced, as for instance the specific heat and magnetization data as well as the continuum of states seen in neutron scattering. Here we investigate the thermodynamic response to magnetic fields applied along the global [110] direction using specific heat measurements and fits using numerical methods, and solve the corresponding magnetic structure using neutron diffraction. Specific heat data in moderate fields are reproduced well, however, at high fields the agreement is not satisfactory. We especially observe a two-step release of entropy, a finding that demands a review of both theory and experiment. We address it within the framework of three possible scenarios, including an analysis of the crystal field Hamiltonian not restricted to the two-dimensional single-ion doublet subspace. We conclusively rule out two of these scenarios and find qualitative agreement with a simple model of field misalignment with respect to the crystalline direction. We discuss the implications of our findings for [111] applied fields and for future experiments on Ce2Hf2O7 and its sister compounds.
@@ I Introduction
In pyrochlores, the magnetic ions form a lattice of tetrahedra connected by their summits, and the basic motivation came from spin ice physics, originally discovered in Ho2Ti2O7 and Dy2Ti2O7 [22, 23, 24, 25]. In these materials, a strong crystal field forces dipole magnetic moments to align along or against the local axis (local axis) directed from the site to the center of the tetrahedron. Such systems do not show any long range order down to mK temperatures, instead their properties are governed by magnetic configurations formed by an organizing principle, the ‘2-in 2-out ice rule’, emerging from effective ferromagnetic interactions. This ice rule constraint manifests itself in various ways, notably the observation of Pauling’s residual entropy [24] and the existence of a classical Coulomb liquid with characteristic ‘pinch points’ [26] accessible in neutron scattering.
@@
One approach to find QSLs has been to look for pyrochlore magnets with relaxed constraint on the Ising anisotropy of the magnetic moments. This approach, however, has proved difficult. For example, the case of Yb2Ti2O7 [27, 28] was anticipated to be a QSL but was found to have a ferromagnetic ground state [29, 30, 31, 32, 33]. Alternatively, a promising route for QSL physics was found with materials based on the non-Kramers trivalent Pr3+ ion having Ising-like dipole magnetic moments [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], and where interactions between electric quadrupoles [45] and residual non-magnetic disorder [41, 43] can both be source of quantum fluctuations to stabilize a QSL.
@@
More recently, the case of trivalent rare-earth ions that can stabilize a ‘dipole-octupole’ ground state doublet, such as Nd or Ce, has been actively investigated and also appears very promising [46, 47, 48, 49, 50]. As in most rare-earth pyrochlores, since the crystal-electric field (CEF) ground state doublet is well separated from the rest of the spectrum by a gap of few tens of meV, microscopic descriptions use an effective spin 1/2 degree of freedom. The special feature of these Kramers dipole-octupole doublets lies in the fact that the component of the pseudo-spin essentially represents the magnetic (dipolar) moment, while the the and components are octupolar in nature. This has important consequences on the nature of the low energy effective description. Following previous studies, the theoretical calculations presented in this work are based on a minimal nearest neighbor (nn) XYZ Hamiltonian incorporating all symmetry-allowed effective interactions between these pseudo-spins on the pyrochlore lattice [51, 52, 53, 54, 55, 56]. Accounting for the applied magnetic field when projected onto the ground doublet subspace gives:
@@
The basis is chosen such that is defined in the local direction on a given rare-earth site and . and are effective coupling strengths and refers to nearest neighbor pairs of sites. We note that, in the chosen basis with , the Zeeman term couples only the local component of the magnetic field to the dipolar component while the octupolar ( and ) components do not contribute to this term.
@@
Depending on the relative strength of , different situations can emerge. For instance, in Nd2Zr2O7, is dominant and negative, leading to an antiferromagnetic all-in–all-out ground state, with however significant dynamics arising from a non-negligible ‘transverse’ term [49, 57, 58, 50]. In the Ce3+-based pyrochlore materials Ce2Sn2O7, Ce2Zr2O7 and Ce2Hf2O7 [48, 59, 46, 47, 60, 56, 61, 62], previous work determined a dominant positive interaction between and/or pseudo-spin components, hence a new class of spin ices where the ice rule is ‘octupolar’. As a result, the magnetic diffuse scattering in Ce2Sn2O7 does not show the familiar characteristic signature observed in the dipolar spin ice systems. Some magnetic diffuse scattering, however, arises at high momentum transfers and is observable with neutrons thanks to the weak coupling of octupoles to neutrons [48, 56]. Furthermore, in all three Ce3+-based pyrochlores, a continuum of spectral weight [46, 47] – the hallmark of fractionalized excitations in a QSL, has been observed in neutron scattering in zero applied magnetic field.
For Ce2Hf2O7, effective exchange parameters were reported previously by us [61] with a dominant interaction:
@@
(all ’s in meV, dimensionless), which locates the material of the present study in the -flux U(1) QSL phase [63, 51, 64, 65, 66, 53, 52], as suggested also for Ce2Zr2O7 [47, 46, 56, 67]. The interaction parameters result in a rich phase diagram for cerium pyrochlores, and their tiny energy scales result in a potentially large response to small changes in the material composition. For instance, this is illustrated in the markedly different exchange parameters in Ce2Zr2O7 and Ce2Hf2O7 despite Zr4+ and Hf4+ sharing very similar ionic radii and electronegativities. The highly tunable nature of these tiny energy scales may also explain the contrasted magnetic neutron scattering response of Ce2Sn2O7 samples prepared at lower temperature [62] compared to other studies [59, 48, 59, 68]. The methodology to deduce reliable sets of exchange parameters from the available experimental data is a central problem in the study of these materials.
@@
In the case of dipolar systems, the application of external magnetic fields has been a valuable resource for building confidence in the microscopic description of such materials [27, 25]. Large fields spin polarize the system allowing direct measurement of the energy- and momentum-resolved spin waves (for example, see the case of Yb2Ti2O7 [27, 30, 31, 28] and references therein). Using linear spin wave theory, it is then possible to accurately determine the underlying interactions in these dipolar spin ices. The case of dipolar-octupolar magnets is starkly different [51, 64, 65, 66, 53, 52]: a field-induced magnetic order can exist, but with no spectral weight associated with the octupolar low energy excitations measured with neutrons at low momentum transfers. Absence of spin waves has also been noted in Ce2Zr2O7, despite the measurement of a field-induced magnetic order in a finite magnetic field [67] These features outline the importance of methods to evaluate exchange interactions of cerium pyrochlores from measured data.
@@
Given this prelude to dipole-octupole pyrochlore magnets, in this work we present neutron diffraction, magnetization and specific heat data on Ce2Hf2O7 for magnetic fields applied along the crystallographic [110] direction. Strikingly, we infer a two-step release of entropy from the specific heat measurements, a finding that demands a review of the experiment and exploration of multiple theoretical scenarios. Using a combination of quantum many-body calculations and analytical insights, we present strong evidence for the importance of a field misalignment away from the [110] axis as the origin of the peculiar evolution of entropy as a function of the applied field strength. We also revisit results for the [111] field direction, previously addressed elsewhere [61].
@@
The rest of the paper is organized as follows. In section II we describe the setup of the pyrochlore lattice, discussing the phases expected for high magnetic fields. In section III we discuss important details of our experiments, and the results of measurements for magnetic fields, highlighting the agreement between new data for moderate [110] fields with the prediction from our previously determined model parameters given in Eq. (2). Section IV is devoted to our theoretical analysis for [110] fields, while Section V provides, for completeness, a similar analysis for the [111] fields, with simple torque calculations confirming that the effects of a misalignment are far less important in that case. Finally, in Section VI we conclude and discuss the implication of our work on future experiments.
@@ II Pyrochlore Setup
We distinguish the global coordinate system, which we denote with , from the local coordinate system at each rare earth site, which is given, following Ref. [51], as follows:
@@
Equation (1) indicates that only the local components are of consequence in determining the coupling of the applied external magnetic field to a given pseudo-spin.
An interesting situation arises for a magnetic field of strength applied perfectly along [110], i.e., and in the global basis. The Zeeman term is exactly zero for sublattices 1 and 2, i.e. the spins do not directly couple to the applied [110] magnetic field, for arbitrary field strengths. The Zeeman term is non-zero for sublattices 0 and 3, i.e., the sites. For the case of large field strengths, the ground state corresponds to the pseudospin on sublattice 0 pointing “out” and the one on sublattice 3 pointing “in” i.e. along the local axes, away and towards the center of the tetrahedron, respectively. This configuration has been qualitatively depicted in Fig. 1(b).
If the spins were truly free, the ground state would be exactly degenerate resulting in a residual entropy of per site. In practice, the spins do couple to the spins which in turn couple to the magnetic field, thus the spins are essentially free only at temperature scales above the scale of effective magnetic interactions. One should, however, expect to see the signatures of this effective decoupling in the specific heat and the associated magnetic entropy, especially at high fields where the dominant scale is set by the Zeeman term. Rather curiously, we find that our high-field experiments do not exactly adhere to this expectation, a finding we address further in the next sections.
@@ III Experimental results
We now present the details of the experiments performed, along with the key outcomes and puzzles they pose. In subsection III.1, we show neutron diffraction data in [110] fields, which provide a picture consistent with the spins ordering while the spins remain disordered. We then discuss in subsection III.2 specific heat results to be compared with numerical calculations presented in section IV.
@@ III.1 Magnetic neutron diffraction and magnetization
The induced magnetic structure under application of a magnetic field parallel to the direction was investigated using single crystal neutron diffraction (Zebra, SINQ, PSI). The single crystal sample of Ce2Hf2O7 was cooled down using a dilution refrigerator inside a 10 T vertical cryomagnet. Reflections in and out of the horizontal scattering plane ( degrees) were collected using thermal neutrons (1.18 Å) via a 3He point detector. The neutron beam was monochromatized and vertically focused using the (311) Bragg reflection of a germanium monochromator. Measurements took place at 1.5 K, 0.2 K and 0.05 K, under the application of a magnetic field = 0 T, 0.5 T, 2 T and 8 T. Magnetic Bragg intensities were extracted by subtraction of their respective nuclear contributions measured without magnetic field. All magnetic reflections could be accounted for using a propagation vector. Other directions of reciprocal space were investigated in order to check for magnetic Bragg scattering of a different propagation vector, which confirmed that the magnetic structure can be described within the chemical unit cell.
Fits of the experimental integrated intensities corrected for Lorentz factor were performed using the FullProf Suite program [69] assuming the formation of and chains and a 2-in-2-out local constraint (the fit is displayed in Fig. 2), owing to the strong anisotropy and frustration experienced by the magnetic moments [53]. Across all temperatures and non-zero magnetic fields, the determined magnetic structures revealed a net magnetic moment on the sites, lying along their respective local easy axis, polarized in the field direction and increasing with field strength. No net magnetic moment on the sites could be identified for = 0 T, 0.5 T, 2 T. Interestingly, for a magnetic field strength of = 8 T, allowing a small magnetic moment (0.100.05 ) parallel to the magnetic field leads to a slight improvement of the fit. This could be due to the increased number of fit parameters (and therefore be artificial) or could alternatively be real, potentially stemming from a slight misalignment of the field with respect to the ideal direction or due to a reduced single-ion anisotropy inherited from defective CEF environments [60].
Magnetization vs field was measured using SQUID magnetometers equipped with a miniature dilution refrigerator developed at the Institut Néel-CNRS Grenoble [70]. We report data for the magnetic field applied along the [110] direction, as shown in Fig. 2(b) for data at 0.08 K and in Fig. 3(d) at four representative temperatures. The magnetization of the sample was also calculated based on the magnetic structure determined from neutron diffraction at 0.05 K and compared with the macroscopic magnetization at 0.08 K (Fig. 2(b)), showing a good agreement with, and further corroborating, the -chain/-chain structure ansatz.
@@ III.2 Specific heat
A thin slice of Ce2Hf2O7 was prepared for specific heat measurement using the single crystal reported in [60] and a tungsten wire saw. The [110] orientation of the slice’s faces was assessed using a x-ray Laue diffractometer. Specific heat measurements between 0.05 K and 0.8 K were performed using a home-built calorimeter inserted in a dilution refrigerator at Université de Sherbrooke. For the rest of the paper, we will refer to this as “experimental setup A”. In this setup, the heater and the thermometer were glued directly to the sample using GE varnish. This assembly was suspended on zylon wires to decouple it thermally from the copper mount. Electrical contacts were made using 7 µm diameter superconducting NbTi wires. The thermal conductivity between the sample assembly and the copper mount was adjusted using a manganin wire, in order to have suitable relaxation times for the measurements. The manganin heat link was 1.7 mm long with a diameter of 0.1 mm. As well as providing the thermal link to the refrigerator, this short length of wire also provided significant structural support to the sample in addition to the very thin zylon filaments. The sample was oriented to apply the magnetic field along the [110] crystallographic direction. Data were acquired using the quasi-adiabatic heat pulse method.
@@
Specific heat measurements between 0.4 K and 15 K were carried out using a Quantum Design physical properties measurement system (PPMS), with vertically applied magnetic fields up to 6 T. This is referred to as “experimental setup B” for the rest of the paper. For this setup, the sample was mounted with its smoothest side facing down towards the puck’s platform using an Apiezon grease, allowing the application of a magnetic field parallel to the [110] crystallographic direction. The platform itself is suspended by four wires, which also provide electrical connections to the heater and thermometer. Data measured on the nonmagnetic analog La2Hf2O7 were used to subtract lattice contributions.
@@
Figures 3(a),(b),(c) and 4 show experimental specific heat data for Ce2Hf2O7 in a magnetic field along the [110] direction in both setup A (green) and B (black), for low ( T) and high field strengths ( T) respectively. Note that for the latter case, only data from setup B is available. The experimental results reveal a prominent two-step release of entropy, which appears as two outstanding peaks in the specific heat – the lower temperature feature being somewhat sharp and the high temperature feature being a broader bump. Note that such a two-step entropy release, if present, is not prominent for fields applied in the [111] direction [61]. The temperature of both peaks in the specific heat increases with the applied field strength.
@@
As emphasized earlier, the field applied along the [110] direction separates the spins in two types of chains, labeled and . Owing to the amplitude of the various effective interactions, the high-temperature bump arises from the Zeeman term of the spins, scaling linearly with the field strength. The low temperature feature indicates correlations on the -chains, as investigated in more detail in the next section.